OpenStudy (anonymous):

Solve –3(2x – 4) = 24 Solve –3(2x – 4) = 24 @Mathematics

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

yes, x=-2

OpenStudy (anonymous):

Answer is not important. How the answer is obtained is more important. -3(2x-4)=24 -3*2x -(-3)*(-4) =24 -6x + 12 = 24 -6x = 24-12 -6x = 12 x = -12/6 x = -2

OpenStudy (anonymous):

not in university

OpenStudy (anonymous):

professor doesn't give a sh!t how u found the answer

OpenStudy (anonymous):

Sorry...I do.